Amounts for Materials

Big Timber Furniture Company manufactures furniture. Big Timber Furniture uses a job order cost system. Balances on June 1 from the materials ledger are as follows:

Fabric $27,700
Polyester filling 8,300
Lumber 62,000
Glue 2,700

The materials purchased during June are summarized from the receiving reports as follows:

Fabric $136,300
Polyester filling 189,500
Lumber 373,500
Glue 12,900

Materials were requisitioned to individual jobs as follows:

Fabric Polyester Filling Lumber Glue Total
Job 101 $51,800 $65,300 $174,600 $291,700
Job 102 39,900 71,000 152,800 263,700
Job 103 36,800 48,200 85,000 170,000
Factory overhead-indirect materials $6,700 6,700
Total $128,500 $184,500 $412,400 $6,700 $732,100

The glue is not a significant cost, so it is treated as indirect materials (factory overhead).

a. Determine the total purchase of materials in June.
$

b. Determine the amounts of materials transferred to Work in Process and Factory Overhead for the requisition of materials in June.

Materials transferred to Work in Process $
Materials transferred to Factory Overhead $

c. Determine the June 30 balances that would be shown in the materials ledger accounts.

Balance, June 30
Fabric $
Polyester Filling $
Lumber $
Glue $
Answer #1
a
Total purchase of materials 712200 =136300+189500+373500+12900
b
Materials transferred to Work in Process 725400 =128500+184500+412400
Materials transferred to Factory Overhead 6700
c
Balance, June 30
Fabric 35500 =27700+136300-128500
Polyester Filling 13300 =8300+189500-184500
Lumber 23100 =62000+373500-412400
Glue 8900 =2700+12900-6700